How long does at take for longitudinal wave to travel distance of 100mm in steel

Grade:12th pass

2 Answers

sree lakshmi
4 years ago
In stainless steel velosoty of longitudinal wave is5664m/s
Time = distance/velocity
     D    = 100mm =0.1m
     Time=0.1/5664
       = 1.76\times 10^{-5} second
Khimraj
4 years ago
 
D    = 100mm =0.1m
     Time=0.1/5664
       = 1.76\times 10^{-5} second
,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,